# NAP-Tuning: Neural Augmented Prompt Tuning for Adversarially Robust Vision-Language Models

### v1

@@ IV-D2 Principled Loss Formulation
AGS Vanilla $46.1$ $85.0$ $29.8$ $11.8$ $13.4$ $50.9$ $49.1$ $73.6$ $47.5$ $49.6$ $46.9$ $45.8$
FAP 52.0 89.9 54.4 $31.6$ $21.7$ 63.4 $81.0$ 80.1 $48.8$ 61.1 $61.5$ $58.7$
AdvPT $48.6$ $88.0$ $48.5$ $29.9$ $23.9$ $50.5$ $71.9$ $74.9$ $57.3$ $57.5$ $60.1$ $55.6$
AdvPT-V $47.7$ $85.9$ $33.6$ $3.6$ $12.3$ $49.3$ $46.7$ $73.6$ $37.9$ $48.9$ $44.3$ $44.0$
AdvPT-VLI $50.6$ $83.8$ $26.4$ $14.7$ $6.7$ $49.3$ $43.0$ $76.1$ $33.9$ $43.7$ $41.8$ $42.7$
AdvPT-VLJ $51.2$ $80.0$ $26.7$ $14.6$ $7.2$ $48.0$ $39.5$ $74.5$ $32.0$ $44.5$ $43.1$ $41.9$
[…]
PGD Vanilla $2.0$ $29.1$ $4.9$ $0.2$ $0.0$ $7.4$ $1.7$ $4.2$ $0.4$ $1.5$ $3.0$ $6.7$
FAP $23.4$ $67.1$ $32.8$ $0.4$ $7.0$ $26.2$ $57.3$ $40.0$ $17.6$ $29.8$ $31.5$ $23.3$
AdvPT $2.8$ $31.7$ $10.0$ $0.4$ $0.8$ $1.8$ $8.7$ $3.2$ $2.0$ $2.9$ $4.5$ $5.2$
AdvPT-V $19.0$ $64.6$ $19.8$ $0.3$ $2.9$ $15.0$ $23.0$ $36.3$ $9.6$ $18.9$ $18.3$ $20.4$
AdvPT-VLI $20.7$ $62.1$ $16.3$ $5.7$ $1.5$ $15.9$ $19.3$ $37.6$ $8.4$ $16.2$ $17.7$ $20.2$
AdvPT-VLJ $20.9$ $59.9$ $15.2$ $10.1$ $2.0$ $15.3$ $18.8$ $37.5$ $9.6$ $16.9$ $18.2$ $18.9$
[…]
@@
FAP $2.0$ $1.3$ $5.0$ $1.6$ $3.5$ $2.0$ $1.9$ $1.5$ $1.0$ $3.2$ $2.3$ $2.2$
AdvPT $0.0$ $0.6$ $0.5$ $0.1$ $0.2$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$ $0.2$
AdvPT-V $8.3$ $44.8$ $12.4$ $0.0$ $0.5$ $5.4$ $8.8$ $16.3$ $2.8$ $7.4$ $7.2$ $10.0$
AdvPT-VLI $9.3$ $44.2$ $10.5$ $0.2$ $0.4$ $5.6$ $7.6$ $16.4$ $2.1$ $6.2$ $6.8$ $10.1$
AdvPT-VLJ $10.1$ $44.7$ $10.3$ $1.2$ $0.6$ $6.1$ $7.9$ $18.9$ $2.3$ $7.2$ $8.3$ $11.0$
@@ V Experiments
This section systematically evaluates the effectiveness of our proposed NAP-Tuning. We first introduce the experimental setup, then validate the key components of our method through a series of carefully designed experiments, and comprehensively compare with existing approaches.
@@ V-A1 Datasets and Models
To comprehensively evaluate the effectiveness of our approach, we conduct tests on 11 widely used image classification datasets, including ImageNet [21], Caltech101 [22], DTD [23], EuroSAT [24], FGVC Aircraft [25], Food101 [26], Oxford Flowers [27], Oxford Pets [28], Stanford Cars [29], SUN397 [30], and UCF101 [31]. These datasets encompass a diverse range of visual tasks from fine-grained recognition to scene classification. We follow the training and testing splits defined in [7]. For the ImageNet test set, consistent with prior adversarial attack studies [32, 33, 34], we randomly sample 1,000 images, ensuring one image per class.
Our experiments are conducted on the CLIP model, where the default configuration employs the publicly available ViT-B/16 architecture unless otherwise specified. We also include results for the ViT-B/32 model as a supplementary evaluation. In accordance with previous work, we use hand-crafted prompts as textual inputs (e.g., “a photo of a <class>, a type of pet” for Oxford Pets).
V-A2 Attack Methods
To evaluate adversarial robustness, we implement both white-box and black-box adversarial attacks with more stringent settings compared to our conference version. For white-box attacks, we use learnable textual vectors to construct cross-entropy loss and employ stronger PGD-100 [4] and AutoAttack [35]. For black-box attacks, we utilize more advanced methods such as CWA [36] and AGS [37].
V-A3 Defense Baselines
We compare our method against several strong baselines, including variants of AdvPT and the recent FAP [5]:
  • AdvPT: the original Adversarial Prompt Tuning method, which performs prompt tuning exclusively on the text side. It adapts the textual input prompts to improve robustness without modifying the visual inputs.
  • AdvPT-V: an extension of AdvPT that applies prompt tuning on the vision side. Specifically, it introduces additional learnable prompts into the visual encoder, enabling adaptation of the visual input space to adversarial perturbations.
  • AdvPT-VLJ (vision-language joint prompt tuning): this variant simultaneously introduces learnable prompts to both the vision and text modalities. Moreover, it establishes a mapping between the vision and language prompts, allowing joint adversarial tuning of both modalities. This structure follows a design similar to MaPLe [8].
  • AdvPT-VLI (vision-language independent prompt tuning): a modification of AdvPT-VLJ that removes the mapping between vision and language prompts. In this setup, the learnable prompts for the visual and textual inputs are optimized independently, without explicit cross-modal interactions.
  • FAP [5]: a recent method proposed for few-shot adversarial robustness, which carefully designs the objective function.
V-A4 Implementation Details
@@
Our training framework spans 90 epochs with a batch size of 512. We adopt the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is modulated through a cosine annealing schedule. For adversarial example generation during training (as defined in Equation (12)), we implement PGD-5 with a default perturbation magnitude $\epsilon=1/255$ for both training and testing phases, though we analyze the effects of different $\epsilon$ values in subsequent sections. For the TokenRefiner $R$, we implement a default 2-layer neural network architecture, with the impact of network depth examined in our ablation studies. Regarding the multi-layer prompt architecture, our default configuration utilizes all 12 transformer layers, with layer-specific effects analyzed in detailed experiments. The hyperparameter $\alpha_{0}$ in Equation (17) is set to 5.0. All experiments are conducted on NVIDIA A800 80GB GPUs.
V-B Main Results
@@
The main evaluation results are summarized in Tables I and II. Overall, NAP-Tuning achieves highly competitive clean accuracy compared to existing baselines. While its clean performance is slightly lower than that of AdvPT, this is expected, as AdvPT explicitly prioritizes clean accuracy in the robustness-accuracy trade-off, often at the cost of adversarial robustness.
In terms of adversarial robustness, NAP-Tuning demonstrates clear and consistent advantages. Under black-box attacks, particularly CWA and AGS, our method outperforms all baselines across both ViT-B16 and ViT-B32 backbones, highlighting its superior generalization to unseen adversarial threats. More importantly, under white-box attacks such as PGD and AutoAttack, NAP-Tuning exhibits significant improvements. AutoAttack is widely regarded as the most rigorous and comprehensive white-box evaluation benchmark, combining several strong attacks to provide a thorough assessment of model robustness. Under this challenging setup, NAP-Tuning achieves substantial gains over all competitors: on ViT-B16, it attains 51.4% on average under AutoAttack, outperforming the strongest baseline (AdvPT-V) by 33.5%; on ViT-B32, it reaches 44.0%, exceeding the best competitor (AdvPT-VLJ) by 33.0%. These results clearly demonstrate the robustness of NAP-Tuning against powerful, diverse adversarial threats.
V-C Multi-layer Prompt Architecture
@@
The results reveal several key insights. First, as shown in the average performance (Fig. 4l), both clean and robust accuracy generally improve as the number of layers increases, confirming the effectiveness of our hierarchical prompting strategy. Interestingly, we observe dataset-specific patterns related to task complexity. For more challenging datasets such as ImageNet (Fig. 4f), SUN397 (Fig. 4j), and Food101 (Fig. 4e), performance follows an inverted U-shape, peaking at intermediate layer depths before declining. This pattern indicates mild overfitting when increasing parameter capacity without corresponding increases in training examples (shot=16). The highlighted regions in these plots mark the optimal layer ranges for these datasets.
These findings reveal the potential for further performance enhancements through larger training sets and dataset-specific layer optimization, suggesting that improved robustness can be achieved through straightforward data augmentation. Nevertheless, our method demonstrates strong performance even with the default configuration (layer=12) as shown in Table I and Table II, consistently outperforming baseline approaches across diverse datasets.
V-D Adversarial Regularization Parameter
We analyze the impact of adversarial regularization weight ($\alpha$) on both standard (clean) and adversarial (robust) accuracy. Fig. LABEL:fig:alpha presents results across $\alpha$ values from 0.01 to 100, with panel (a) showing both metrics on a single y-axis and panel (b) using dual y-axes to highlight relative trends.
The results show that clean accuracy remains stable (98.5%-99.1%) across all $\alpha$ values, while robust accuracy increases substantially from 79.9% at $\alpha$=0.01 to 93.9% at $\alpha$=2.0, with diminishing gains thereafter. For $\alpha\geq 5.0$, robust accuracy consistently exceeds 94%. Rather than specifically optimizing for the peak performance at $\alpha$=20.0 (94.4% robust accuracy), we adopt $\alpha$=5.0 as our representative configuration, which achieves 98.8% clean accuracy and 94.2% robust accuracy. These findings demonstrate that our approach effectively balances clean accuracy and adversarial robustness across a broad range of $\alpha$ values, exhibiting strong performance without requiring fine-tuned parameter selection.
V-E TokenRefiner Architecture
We analyze the impact of TokenRefiner network depth on model performance using Oxford Flowers (simpler task) and ImageNet (more challenging task). Fig. LABEL:fig:token presents clean and robust accuracy across different TokenRefiner layer configurations. Results demonstrate that TokenRefiner is essential for model convergence—its absence (layer = 0) prevents the model from converging under 5-step PGD training, resulting in poor performance (4.6% clean accuracy on ImageNet, 25.8% on Oxford Flowers). Performance improves substantially as layers increase from 0 to 2, with ImageNet clean accuracy rising from 4.6% to 54.4% and robust accuracy from 1.9% to 28.0%. Similarly, Oxford Flowers shows improvements from 25.8% to 96.1% (clean) and from 15.1% to 86.6% (robust). Performance stabilizes beyond 2 layers, leading us to select a 2-layer architecture for our final implementation. Notably, unlike previous approaches such as FAP [5] and APD [6], which lack specialized prompt structure designs and are therefore limited to weaker adversarial search strategies, our TokenRefiner enables convergence under stronger adversarial perturbations. The consistent pattern across datasets of varying complexity suggests our approach generalizes well without requiring task-specific architectural adjustments.
V-F Perturbation Magnitude
We analyze the impact of perturbation budgets during both training and testing phases. Table III shows results on the Flowers dataset with varying $\epsilon$ values. The results reveal a clear trade-off between clean accuracy and adversarial robustness. As training perturbation budget increases from $\epsilon=1$ to $\epsilon=4$, clean accuracy decreases from 96.1% to 87.5%, while robustness against stronger attacks improves substantially. Models trained with larger $\epsilon$ values perform significantly better when tested under matching or higher perturbation budgets—a model trained with $\epsilon=4$ achieves 55.9% accuracy when tested at the same level, versus only 17.1% for a model trained with $\epsilon=1$. Models experience rapid performance degradation when tested with perturbations larger than their training budget. This observation suggests that, similar to traditional adversarial training, adversarial prompt tuning exhibits a clear trade-off between clean accuracy and adversarial robustness.

### v2

@@ IV-D2 Principled Loss Formulation
M-Attack Vanilla $18.3$ $33.7$ $15.0$ $11.9$ $13.1$ $31.5$ $18.7$ $35.2$ $19.5$ $28.4$ $22.6$ $22.5$
FAP $39.7$ $65.6$ $28.3$ $9.6$ $21.5$ $48.4$ $39.8$ $51.5$ $48.8$ $40.2$ $32.0$ $38.7$
APD $38.4$ $68.2$ $22.9$ $16.1$ $16.9$ $40.6$ $33.7$ $53.6$ $39.6$ $36.8$ $30.1$ $36.1$
AdvPT $32.0$ $41.5$ $11.9$ $9.2$ $6.7$ $33.7$ $25.8$ $43.1$ $29.8$ $27.6$ $42.1$ $27.6$
AdvPT-V $40.9$ $69.3$ $33.6$ $3.6$ $12.3$ $49.3$ $46.7$ $50.9$ $37.9$ $48.9$ $44.3$ $39.8$
AdvPT-VLI $42.6$ $65.8$ $28.3$ $11.9$ $13.1$ $42.7$ $34.8$ $52.5$ $47.6$ $46.0$ $42.1$ $38.9$
AdvPT-VLJ $42.9$ $66.1$ $28.2$ $14.8$ $9.5$ $45.6$ $33.9$ $55.8$ $48.1$ $44.7$ $45.3$ $39.5$
NAP-Tuning 44.6 76.5 37.1 51.9 36.8 52.1 67.4 57.3 67.7 50.9 48.3 53.7
PGD Vanilla $2.0$ $29.1$ $4.9$ $0.2$ $0.0$ $7.4$ $1.7$ $4.2$ $0.4$ $1.5$ $3.0$ $4.9$
FAP $24.5$ $66.8$ $27.1$ $28.3$ $19.2$ $19.8$ $69.7$ $40.5$ $29.9$ $30.3$ $36.9$ $35.7$
APD $18.4$ $66.0$ $30.8$ $22.4$ $20.2$ $26.1$ $64.5$ $33.9$ $36.8$ $30.4$ $36.8$ $35.1$
AdvPT $2.8$ $31.7$ $10.0$ $0.4$ $0.8$ $1.8$ $8.7$ $3.2$ $2.0$ $2.9$ $4.5$ $6.3$
AdvPT-V $19.0$ $64.6$ $19.8$ $0.3$ $2.9$ $15.0$ $23.0$ $36.3$ $9.6$ $18.9$ $18.3$ $20.7$
AdvPT-VLI $20.7$ $62.1$ $16.3$ $5.7$ $1.5$ $15.9$ $19.3$ $37.6$ $8.4$ $16.2$ $17.7$ $20.1$
AdvPT-VLJ $20.9$ $59.9$ $15.2$ $10.1$ $2.0$ $15.3$ $18.8$ $37.5$ $9.6$ $16.9$ $18.2$ $20.4$
NAP-Tuning 26.5 77.9 34.2 49.1 25.4 28.4 82.4 41.3 52.4 31.5 45.0 44.9
@@
FAP $10.9$ $42.8$ $8.5$ $1.8$ $1.0$ $7.1$ $16.7$ $26.2$ $2.4$ $10.0$ $12.5$ $12.7$
APD $9.0$ $46.5$ $10.1$ $0.2$ $0.5$ $6.5$ $14.8$ $17.9$ $1.9$ $9.4$ $12.2$ $11.7$
AdvPT $0.0$ $0.6$ $0.5$ $0.1$ $0.2$ $0.0$ $0.0$ $0.1$ $0.0$ $0.1$ $0.0$ $0.1$
AdvPT-V $8.3$ $44.8$ $12.4$ $0.0$ $0.5$ $5.4$ $8.8$ $16.3$ $2.8$ $7.4$ $7.2$ $10.4$
AdvPT-VLI $9.3$ $44.2$ $10.5$ $0.2$ $0.4$ $5.6$ $7.6$ $16.4$ $2.1$ $6.2$ $6.8$ $9.9$
AdvPT-VLJ $10.1$ $44.7$ $10.3$ $1.2$ $0.6$ $6.1$ $7.9$ $18.9$ $2.3$ $7.2$ $8.3$ $10.7$
@@ V-A1 Datasets and Models
To comprehensively evaluate the effectiveness of our approach, we conduct tests on 11 widely used image classification datasets, including ImageNet [24], Caltech101 [5], DTD [2], EuroSAT [6], FGVC Aircraft [17], Food101 [1], Oxford Flowers [21], Oxford Pets [22], Stanford Cars [9], SUN397 [32], and UCF101 [27]. These datasets encompass a diverse range of visual tasks from fine-grained recognition to scene classification. We follow the training and testing splits defined in [40]. For the ImageNet test set, consistent with prior adversarial attack studies [4, 29, 33], we randomly sample 1,000 images, ensuring one image per class. Our experiments are conducted on the CLIP model. In accordance with previous work, we use hand-crafted prompts as textual inputs (e.g., “a photo of a <class>, a type of pet” for Pets).
To evaluate adversarial robustness, we implement both white-box and black-box adversarial attacks. For white-box attacks, we use PGD-100 [16] and AutoAttack [3]. Consistent with our training protocol, we use a perturbation magnitude of $\epsilon=1/255$. This provides a direct evaluation of the robustness learned at the training budget. For black-box attacks, we utilize the strong M-Attack [11]. As black-box attacks are inherently less effective, and to ensure a challenging and meaningful evaluation, we follow standard evaluation practice and use a larger perturbation budget of $\epsilon=8/255$.
V-A2 Defense Baselines
We compare NAP-Tuning against state-of-the-art prompt-based defenses. These include the original AdvPT (text-only) and its multi-modal extensions: AdvPT-V (vision-only), AdvPT-VLI (independent dual-modal), and AdvPT-VLJ (joint mapping akin to MaPLe [8]). We also include recent strong baselines: FAP [41], which optimizes a few-shot robustness objective, and APD [15], which utilizes bimodal distillation from a clean teacher.
V-A3 Implementation Details
@@
Our training framework spans 90 epochs with a batch size of 32. We adopt the AdamW optimizer with an initial learning rate of $1\times 10^{-4}$, which is modulated through a cosine annealing schedule. For adversarial example generation during training, we implement PGD-5 with a default perturbation magnitude $\epsilon=1/255$ for training phases, though we analyze the effects of different $\epsilon$ values in subsequent sections. For the TokenRefiner $R$, we implement a default 2-layer neural network architecture, with the impact of network depth examined in our ablation studies. Regarding the multi-layer prompt architecture, our default configuration utilizes all 12 transformer layers, with layer-specific effects analyzed in detailed experiments. The hyperparameter $\alpha_{0}$ in Equation (17) is set to 5.0. All experiments are conducted on a single NVIDIA A800 80GB GPU.
V-B Main Results and Comparisons
We begin by evaluating the overall performance of NAP-Tuning against existing methods. We first compare it to state-of-the-art prompt-based defenses to establish its effectiveness, and then benchmark it against fine-tuning based approaches to highlight the architectural trade-offs.
V-B1 Comparison with State-of-the-Art Baselines
@@
The main evaluation results are summarized in Tables II and III. Overall, NAP-Tuning achieves highly competitive clean accuracy compared to existing baselines. While its clean performance is slightly lower than that of AdvPT, this is expected, as AdvPT explicitly prioritizes clean accuracy in the robustness-accuracy trade-off, often at the cost of adversarial robustness.
NAP-Tuning demonstrates consistent advantages across scenarios. Under the black-box M-Attack, it shows superior generalization, surpassing the strongest baselines by 16.4% on ViT-B16 and 13.9% on ViT-B32. This lead is further amplified under strenuous white-box attacks (e.g., PGD, AutoAttack). Notably, on the rigorous AutoAttack benchmark, our method achieves 48.9% on ViT-B16 and 44.0% on ViT-B32 (see Tables II and III), significantly outperforming all competitors.
V-B2 Comparison with Fine-tuning Defense Methods
We compare NAP-Tuning with TeCoA [18], which fine-tunes CLIP’s encoder layers for adversarial robustness. This comparison directly addresses the fundamental question: should we modify the pre-trained encoder or keep it frozen? Both methods are evaluated on ViT-B32 with adversarial examples generated using PGD-100.
Figure 4: Performance comparison between NAP-Tuning (frozen image encoder) and TeCoA (fine-tuned image encoder).
Impact of Fine-tuning. Fig. 4 reveals a critical trade-off. While TeCoA achieves a smaller robustness gap (15.3% vs. 24.2%), it suffers a catastrophic 32.3% drop in clean accuracy (69.1% $\to$ 36.8%) compared to NAP-Tuning. This failure, consistent across datasets like EuroSAT, confirms that fine-tuning the backbone disrupts the pre-trained multi-modal alignment essential for zero-shot capabilities. In contrast, NAP-Tuning preserves the encoder’s generalization ability by using the Neural Augmentor for targeted feature purification, thereby maintaining strong performance on both clean and adversarial data without compromising the pre-trained representations.
TABLE IV: Training time comparison on Pets dataset. All methods trained for the same number of epochs with identical batch size on a single NVIDIA A800 GPU.
Method Training Efficiency
Time (minutes) Relative
Input-level Prompt Methods
[…]
Deep Prompt Methods
[…]
Backbone Fine-tuning Methods
[…]
V-C Component Ablation Studies
Having demonstrated NAP-Tuning’s superior performance, we now conduct a series of ablation studies to deconstruct our framework. We systematically evaluate the impact of each core design choice, from the high-level prompt architecture down to the specific configuration of our purification module.
V-C1 Effect of Multi-layer Prompt Architecture
@@
Fig. 5 confirms that increasing prompt depth generally enhances both clean and robust accuracy. However, complex datasets (e.g., ImageNet, Food101) exhibit an inverted U-shape trajectory, peaking at intermediate depths. This pattern likely stems from mild overfitting given the limited training samples (16-shot). Despite these task-specific variations, the full-depth configuration (layer=12) yields the superior average performance and consistently outperforms baselines, justifying its adoption as the default setting.
V-C2 Effect of the Neural Augmentor Module
Figure 6: Ablation analysis comparing models with and without the Neural Augmentor (NA) across varying prompt depths. The complete model consistently outperforms the ablated variant.
Complementing the multi-layer analysis, Fig. 6 isolates the specific contribution of the Neural Augmentor. The results unequivocally show that the complete model yields consistently higher robust accuracy across all layer configurations on both ImageNet and Oxford Flowers. A particularly interesting observation is the interaction between module capacity and task complexity. On ImageNet, the complete model peaks earlier (around Layer 4) compared to the ablated variant, which plateaus much later (around Layer 9). This shift suggests that the Neural Augmentor accelerates adaptation by providing necessary structural capacity. While this added capacity introduces a mild overfitting tendency in few-shot settings (manifesting as the inverted U-shape), the substantial performance gain confirms that the benefits of explicit feature purification far outweigh these limitations.
V-C3 Effect of TokenRefiner Modality
Figure 7: Modality-specific TokenRefiner ablation study under PGD-100 attacks. While image-side TokenRefiner provides substantial robustness benefits, the dual-modal approach consistently outperforms single-modality configurations across all datasets.
We analyze the modality-specific contributions in Fig. 7. A clear hierarchy emerges: the Image-only TokenRefiner serves as the primary defense backbone, driving the vast majority of robustness gains. This aligns with the theoretical expectation that direct visual purification is essential against image-based attacks. Interestingly, the Text-only variant—despite operating on unperturbed inputs—still yields considerable robustness improvements. This suggests a mechanism of cooperative feature alignment, where the module dynamically adapts clean text embeddings to better match the purified visual features. Consequently, the full Dual-modal framework consistently outperforms single-branch configurations, confirming that text-side alignment is a critical, non-redundant complement to visual purification.
V-C4 Effect of TokenRefiner Architecture Depth
We investigate the impact of TokenRefiner network depth on model performance using Oxford Flowers (simpler task) and ImageNet (complex task) in Fig. 9(a). The results reveal that the TokenRefiner is not merely an enhancement but a prerequisite for optimization stability. Specifically, removing the module (layer = 0) leads to model collapse under 5-step PGD adversarial training, resulting in near-random performance on both datasets. However, increasing the depth to just 2 layers resolves this instability, yielding dramatic improvements that restore both clean and robust accuracy to competitive levels. Beyond this point, performance saturates, justifying our selection of a 2-layer architecture as the optimal configuration. Notably, this structural stability distinguishes our approach from previous methods like FAP [41] and APD [15]; whereas those methods suffer from convergence issues under strong adversarial search, our TokenRefiner enables robust optimization even under rigorous perturbation constraints. This consistent behavior across datasets of varying complexity confirms that our design generalizes well without necessitating task-specific tuning.
(a) TokenRefiner performance across different network depths. Performance improves dramatically from 0 to 2 layers before stabilizing, with consistent patterns across datasets of varying complexity.
V-D In-depth Analysis
Beyond ablating components, we now provide a deeper analysis of our method’s core mechanism. We investigate how NAP-Tuning works by visualizing its effect on feature space, and explore its compatibility with other defense paradigms.
V-D1 Analysis of Feature Purification
To investigate the effectiveness of different defensive strategies in preserving feature integrity under adversarial attack, we analyze both quantitative distortion measurements and qualitative attention visualizations.
TABLE V: Feature distortion magnitude across different defensive approaches measured using Equation (9). Lower value is better.
Method Feature Distortion ($L_{2}$ Norm)
Vanilla 8.67
AdvPT 8.63
APD 5.01
AdvPT-V 4.87
AdvPT-VLI 4.76
FAP 3.58
NAP-Tuning 1.96